# Invisible neutron decay and light BSM particles

### v1

@@ II.5 A caveat
No search for such three- and four-body final states exist. However, Super-Kamiokande has a search for $n\to\pi^{0}\bar{\nu}$ and $p\to\pi^{+}\bar{\nu}$. The current bounds are Super-Kamiokande:2013rwg :
@@
For the four-body decays of Fig. 1, we can estimate the partial half-live, using results from Heeck:2019kgr , with appropriate replacements:
@@
Here the matrix element is taken to be $\left|W_{0}(\pi)\right|=\left|\langle\pi^{0}|(ud)_{R}d_{R}|n\rangle\right|=0.134\text{GeV}^{2}$ Aoki:2017puj . And again, we assume the Wilson coefficient is equal to one.
@@ III High-energy completions
For 6-fermion operators there are only two possible topologies at tree-level Bonnet:2012kh , using only renormalizable interactions. They are shown in Fig. 2. Note that here the arrows indicate the flow of the particle number. Internal particles can be either scalars, fermions or vectors. For Topology A there are exactly three possibilities to distribute the fermions of ${\cal O}_{1}^{N}$ to the outer legs of the diagram. We list these in Tab. 1.
@@ IV LHC phenomenology
Let us then concentrate on the coloured fermions. Here, we follow largely the discussion in Cepedello:2024qmq . All three coloured fermions can have interactions with SM quarks and the Higgs:
@@
Both ATLAS and CMS have searched for such vector-like quarks in a number of publications. For example, ATLAS has searched for pair-produced VLQs decaying to 1st generation quarks in ATLAS:2024zlo . Limits are given as function of the VLQ mass, for different branching ratios Br($Q\to W+j$) versus Br($Q\to H+j$), assuming Br($Q\to W+j$)+Br($Q\to H+j$)+Br($Q\to Z+j$)=1=1 ATLAS:2024zlo . Derived limits are in the range (9001500)(900-1500) GeV for Br($Q\to W+j$) in the range (10-100) %. In ATLAS:2024gyc ATLAS presented the results of a search for pair-produced VLQs decaying to third generation quarks. Limits range from (11.7)(1-1.7) TeV depending mostly on the branching ratio Br($Q\to W+t/b$). These limits are slightly stronger than the 1st generation quark limits, mostly due to lower backgrounds.
There are also searches for singly produced VLQs. However, in the search ATLAS:2022ozf ATLAS assumes that the VLQ decays to the final state $t+h$ or $W+b$, where the top/bottom is tagged in order to reduce backgrounds. Similarly CMS CMS:2024qdd searched for single VLQ production in the final state $t+h$ and $t+Z$. For $\kappa_{T}=1$, where $\kappa_{T}$ measures the strength of the interaction of the VLQ with electro-weak gauge bosons relative to standard EW couplings, a lower limit on VLQs decaying to top quarks of around $m_{\rm VLQ}\geq 1.9$ TeV is found in ATLAS:2022ozf . There are no searches yet published for singly produced VLQs decaying to second or 1st generation quarks. However, due to much larger backgrounds in such searches, one can expects limits would be considerably weaker.
However, we need to stress that all the VLQ searches discussed up to now depend on the presence of mixing. The neutron decay diagrams, that we study here, however, do not require any of the Yukawa coupling in Eq. (24) to be non-zero. Instead, couplings of the form $F$-$q$-$S$, where $F$, $q$ and $S$ stand for a VLQ, a SM quark and one of the BSM scalars must be present. If the VLQs are pair produced at the LHC they can then decay to a final states containing $q+N_{R}+N_{R}$, i.e. jet plus missing energy. This signal is equivalent to the one used in the standard squark search at the LHC. For example, ATLAS ATLAS:2020syg has searched for pair produced squarks with ${\cal L}=139/$fb. ATLAS derives limits for one generation of squarks (roughly $m_{\tilde{q}}\geq 1.2$ TeV) and 8 degenerate squarks (roughly $m_{\tilde{q}}\geq 1.8$ TeV) for light neutralinos. Since VLQs (being fermions) have cross sections at least $\sim$ 4 times larger than a colour triplet scalar of the same mass, this SUSY search ATLAS:2020syg could be reinterpreted to yield a lower limit on the VLQ mass. We estimate this limit to be roughly $m_{\rm VLQ}\geq 1.5$ TeV, for any of the three coloured VLQs under consideration.
@@
Single or resonant production of LQs requires large Yukawa couplings to give sizeable cross sections. ATLAS searched for resonant LQs ATLAS:2025upm via a lepton-jet signature, with either one or two leptons in the final state. For $Y_{de}=1.0$ LQs with masses below $m_{LQ}=3.4$ TeV are excluded by this search. CMS searched for single LQs in t-channel diagrams CMS:2025iix , leading to di-lepton final states. Both scalar and vector LQs were considered. For scalar [vector] LQs with masses between (15)(1-5) TeV Yukawa couplings larger than (0.31.0)(0.3–1.0) [(0.11.4)(0.1-1.4)] have been excluded.
@@
Finally, diquarks are stringently constrained from dijet searches, since the production is s-channel enhanced. For example, CMS CMS:2019gwf gives a lower limit on the mass of a colour triplet scalar of $m_{DQ}\simeq 7.5$ TeV assuming a Yukawa coupling to two quarks of electromagnetic strength, $e$. The limits derived by CMS are strictly speaking valid only for a scalar diquark coupling to both, up-type and down-type quarks with the same strength. Thus, they apply only to $S_{3,1,-1/3}$.
The searches discussed above are based on luminosities up to 140/fb. Moderate improvements can be expected from the high-luminosity LHC. However, the future FCC-hh FCC:2018vvp would considerably improve sensitivities. For example, FCC:2018byv quotes sensitivities up to $m_{LQ}\sim 8$ ($15$) TeV for pair (singly) produced leptoquarks, decaying to charged leptons. Many other search channels, discussed above, should improve by similar factors.
We will close this section with a short discussion of the $d=12$ operator given in Eq. (1). In the beginning of this section we have argued that a Wilson coefficient of $c_{W}=1$ is often unrealistic in UV models, thus EFT tends to overestimate the masses of the BSM states. This is certainly true also for a $d=12$ operator. [^7] Note that in the decomposition of Eq. (1) up to seven different Yukawa couplings appear. We have used an automated code for operator decomposition, based on the diagrammatic method and described in Cepedello:2022pyx ; Cepedello:2023yao , to decompose Eq. (1) at tree-level. At such a large dimension, there is a proliferation of models. We count 39 topologies, 150 different diagrams and 12713 model variants.
@@ V Summary
In this paper we have studied invisible neutron decay. Limits on nucleon decay modes with three charged leptons Super-Kamiokande:2020tor are much stronger than those of invisible neutron decay ParticleDataGroup:2024cfk . Invisible neutron decay with observable rates can therefore be generated only by operators that do not also generate charged lepton final states. In SMEFT one has to go up to $d=12$ to find such an operator. Current limits on invisible neutron decay correspond then to scales of order $\Lambda\simeq 13$ TeV. An observation of invisible neutron decay in the next round of experiments might then be accompanied by new physics at the future FCC-hh FCC:2018byv .
@@
We have also discussed constraints on the allowed rate for invisible neutron decay, that could be placed by a search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$. Hyper-Kamiokande Hyper-Kamiokande:2018ofw should be able to improve on the limit of $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ from Super-Kamiokande Super-Kamiokande:2013rwg by a considerable factor. We note that the Super-Kamiokande analysis Super-Kamiokande:2013rwg gives only a limit on a two-body final state, but the same type of search can also be used to constrain decays with more than one invisible particle in the final state. Non-observation of $n\to\pi^{0}+3\bar{N}$ at Hyper-Kamiokande might rule out the possibility of observing invisible neutron decay in the foreseeable future.
However, to end this discussion on a more positive note, we mention that the sensitivity of Hyper-Kamiokande and the future JUNO and (the proposed) THEIA experiment to the scales of the invisible neutron decay operators will be rather similar. Observation of invisible neutron decay in JUNO should lead to an excess in the Hyper-Kamiokande search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ and vice versa. The question whether invisible neutron decay or the decay $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ would be discovered first, however, is not so easy to answer. While the future Hyper-Kamiokande experiment Hyper-Kamiokande:2018ofw is much larger than Super-Kamiokande, and thus allows to probe much larger half-lives, already the Super-Kamiokande limit Super-Kamiokande:2013rwg is background dominated. To claim a discovery in this case requires both hundreds of signal events and a detailed understanding of backgrounds. JUNO’s search for the invisible neutron decay JUNO:2024pur , on the other hand, is based on a characteristic triple coincidence arising from the invisible decays of s-shell neutrons in 12C, which leaves very few background events, making a discovery with just a handful of events possible.

### v2

@@ II.5 A caveat
No search for such three- and four-body final states exist. However, Super-Kamiokande has a search for $n\to\pi^{0}\bar{\nu}$ and $p\to\pi^{+}\bar{\nu}$. The current bounds are Abe and others (2014):
@@
For the four-body decays of Fig. 1, we can estimate the partial half-life, using results from Heeck and Takhistov (2020), with appropriate replacements:
@@
Here the matrix element is taken to be $\left|W_{0}(\pi)\right|=\left|\langle\pi^{0}|(ud)_{R}d_{R}|n\rangle\right|=0.134\text{GeV}^{2}$ Aoki et al. (2017). And again, we assume the Wilson coefficient is equal to one.
@@ III High-energy completions
For 6-fermion operators there are only two possible topologies at tree-level Bonnet et al. (2013), using only renormalizable interactions. They are shown in Fig. 2. Note that here the arrows indicate the flow of the particle number. Internal particles can be either scalars, fermions or vectors. For Topology A there are exactly three possibilities to distribute the fermions of ${\cal O}_{1}^{N}$ to the outer legs of the diagram. We list these in Tab. 1.
@@ IV LHC phenomenology
Let us then concentrate on the coloured fermions. Here, we follow largely the discussion in Cepedello et al. (2024). All three coloured fermions can have interactions with SM quarks and the Higgs:
@@
Both ATLAS and CMS have searched for such vector-like quarks in a number of publications. For example, ATLAS has searched for pair-produced VLQs decaying to 1st generation quarks in Aad and others (2024b). Limits are given as function of the VLQ mass, for different branching ratios Br($Q\to W+j$) versus Br($Q\to H+j$), assuming Br($Q\to W+j$)+Br($Q\to H+j$)+Br($Q\to Z+j$)=1=1 Aad and others (2024b). Derived limits are in the range (9001500)(900-1500) GeV for Br($Q\to W+j$) in the range (10-100) %. In Aad and others (2024c) ATLAS presented the results of a search for pair-produced VLQs decaying to third generation quarks. Limits range from (11.7)(1-1.7) TeV depending mostly on the branching ratio Br($Q\to W+t/b$). These limits are slightly stronger than the 1st generation quark limits, mostly due to lower backgrounds.
There are also searches for singly produced VLQs. However, in the search Aad and others (2022) ATLAS assumes that the VLQ decays to the final state $t+h$ or $W+b$, where the top/bottom is tagged in order to reduce backgrounds. Similarly CMS Hayrapetyan and others (2024) searched for single VLQ production in the final state $t+h$ and $t+Z$. For $\kappa_{T}=1$, where $\kappa_{T}$ measures the strength of the interaction of the VLQ with electro-weak gauge bosons relative to standard EW couplings, a lower limit on VLQs decaying to top quarks of around $m_{\rm VLQ}\geq 1.9$ TeV is found in Aad and others (2022). There are no searches yet published for singly produced VLQs decaying to second or 1st generation quarks. However, due to much larger backgrounds in such searches, one can expects limits would be considerably weaker.
However, we need to stress that all the VLQ searches discussed up to now depend on the presence of mixing. The neutron decay diagrams, that we study here, however, do not require any of the Yukawa coupling in Eq. (24) to be non-zero. Instead, couplings of the form $F$-$q$-$S$, where $F$, $q$ and $S$ stand for a VLQ, a SM quark and one of the BSM scalars must be present. If the VLQs are pair produced at the LHC they can then decay to a final states containing $q+N_{R}+N_{R}$, i.e. jet plus missing energy. This signal is equivalent to the one used in the standard squark search at the LHC. For example, ATLAS Aad and others (2021) has searched for pair produced squarks with ${\cal L}=139/$fb. ATLAS derives limits for one generation of squarks (roughly $m_{\tilde{q}}\geq 1.2$ TeV) and 8 degenerate squarks (roughly $m_{\tilde{q}}\geq 1.8$ TeV) for light neutralinos. Since VLQs (being fermions) have cross sections at least $\sim$ 4 times larger than a colour triplet scalar of the same mass, this SUSY search Aad and others (2021) could be reinterpreted to yield a lower limit on the VLQ mass. We estimate this limit to be roughly $m_{\rm VLQ}\geq 1.5$ TeV, for any of the three coloured VLQs under consideration.
@@
Single or resonant production of LQs requires large Yukawa couplings to give sizeable cross sections. ATLAS searched for resonant LQs Aad and others (2025) via a lepton-jet signature, with either one or two leptons in the final state. For $Y_{de}=1.0$ LQs with masses below $m_{LQ}=3.4$ TeV are excluded by this search. CMS searched for single LQs in t-channel diagrams Hayrapetyan and others (2025), leading to di-lepton final states. Both scalar and vector LQs were considered. For scalar [vector] LQs with masses between (15)(1-5) TeV Yukawa couplings larger than (0.31.0)(0.3–1.0) [(0.11.4)(0.1-1.4)] have been excluded.
@@
Finally, diquarks are stringently constrained from dijet searches, since the production is s-channel enhanced. For example, CMS Sirunyan and others (2020a) gives a lower limit on the mass of a colour triplet scalar of $m_{DQ}\simeq 7.5$ TeV assuming a Yukawa coupling to two quarks of electromagnetic strength, $e$. The limits derived by CMS are strictly speaking valid only for a scalar diquark coupling to both, up-type and down-type quarks with the same strength. Thus, they apply only to $S_{3,1,-1/3}$.
The searches discussed above are based on luminosities up to 140/fb. Moderate improvements can be expected from the high-luminosity LHC. However, the future FCC-hh Abada and others (2019b) would considerably improve sensitivities. For example, Abada and others (2019a) quotes sensitivities up to $m_{LQ}\sim 8$ ($15$) TeV for pair (singly) produced leptoquarks, decaying to charged leptons. Many other search channels, discussed above, should improve by similar factors.
We will close this section with a short discussion of the $d=12$ operator given in Eq. (1). In the beginning of this section we have argued that a Wilson coefficient of $c_{W}=1$ is often unrealistic in UV models, thus EFT tends to overestimate the masses of the BSM states. This is certainly true also for a $d=12$ operator. [^8] Note that in the decomposition of Eq. (1) up to seven different Yukawa couplings appear. We have used an automated code for operator decomposition, based on the diagrammatic method and described in Cepedello et al. (2022, 2023), to decompose Eq. (1) at tree-level. At such a large dimension, there is a proliferation of models. We count 39 topologies, 150 different diagrams and 12713 model variants.
@@ V Summary
In this paper we have studied invisible neutron decay. Limits on nucleon decay modes with three charged leptons Tanaka and others (2020) are much stronger than those of invisible neutron decay Navas and others (2024). Invisible neutron decay with observable rates can therefore be generated only by operators that do not also generate charged lepton final states. In SMEFT one has to go up to $d=12$ to find such an operator. Current limits on invisible neutron decay correspond then to scales of order $\Lambda\simeq 13$ TeV. An observation of invisible neutron decay in the next round of experiments might then be accompanied by new physics at the future FCC-hh Abada and others (2019a).
@@
We have also discussed constraints on the allowed rate for invisible neutron decay, that could be placed by a search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$. Hyper-Kamiokande Abe and others (2018) should be able to improve on the limit of $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ from Super-Kamiokande Abe and others (2014) by a considerable factor. We note that the Super-Kamiokande analysis Abe and others (2014) gives only a limit on a two-body final state, but the same type of search can also be used to constrain decays with more than one invisible particle in the final state. Non-observation of $n\to\pi^{0}+3\bar{N}$ at Hyper-Kamiokande might rule out the possibility of observing invisible neutron decay in the foreseeable future.
However, to end this discussion on a more positive note, we mention that the sensitivity of Hyper-Kamiokande and the future JUNO and (the proposed) THEIA experiment to the scales of the invisible neutron decay operators will be rather similar. Observation of invisible neutron decay in JUNO should lead to an excess in the Hyper-Kamiokande search for $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ and vice versa. The question whether invisible neutron decay or the decay $n\to\pi^{0}+E\hskip-6.00006pt/\;\>$ would be discovered first, however, is not so easy to answer. While the future Hyper-Kamiokande experiment Abe and others (2018) is much larger than Super-Kamiokande, and thus allows to probe much larger half-lives, already the Super-Kamiokande limit Abe and others (2014) is background dominated. To claim a discovery in this case requires both hundreds of signal events and a detailed understanding of backgrounds. JUNO’s search for the invisible neutron decay Abusleme and others (2025), on the other hand, is based on a characteristic triple coincidence arising from the invisible decays of s-shell neutrons in 12C, which leaves very few background events, making a discovery with just a handful of events possible.
@@ Appendix A List of the high-energy completions
Appendix B A “black box theorem” for $\Delta(L)=3$ operators
In the introduction, we have argued that with only the particle content of the standard model it is highly unlikely that invisible neutron decay can exist with measurable rates, without also having nucleon decays with charged leptons in the final state with at least comparable rates. In this appendix, we will discuss the basis of this claim in more detail.
As we have mentioned in the introduction, it is possible to forbid all $d=9$ operators by some symmetry. However, as we will now proceed to show, this is not possible at $d=10$. Concretely, we will show that if the $d=12$ SMEFT operator of Eq. (1) exists, than also the $d=10$ $\Delta(L)=3$ operator, ${\cal O}_{d=10}\propto d^{c}d^{c}d^{c}LLLH$, must have a non-zero Wilson coefficient. This proof is modeled on the “black box theorem” for neutrinoless double beta decay Schechter and Valle (1982); Hirsch et al. (2006). This “theorem” is a proof by contradiction. Given two particular operators, one assumes that one of them exists, while the other is forbidden by some symmetry. One then proceeds to show that no such symmetry can exist and that either both operators or none of them will be present in the Lagrangian.
For the formulation of this proof, in addition to the two operators of interest, ${\cal O}_{d=10}$ and ${\cal O}_{d=12}^{\text{W}}$ (see Eq. (1)), we will need the following standard model Yukawa interactions for up and down quarks:
${\cal L}^{\rm Yuk}=Y_{d}\overline{Q_{L}}\cdot Hd_{R}+Y_{u}\overline{Q_{L}}\cdot H^{\dagger}u_{R}\;{\rm+h.c.}$ (32)
Then, assume there exists a global symmetry, which we denote as $G_{\eta}$, under which the SM fields transform as:
$H\stackrel{{\scriptstyle G_{\eta}}}{{\longrightarrow}}\eta_{H}\cdot H$ (33)
$L\stackrel{{\scriptstyle G_{\eta}}}{{\longrightarrow}}\eta_{L}\cdot L,$ $Q_{L}\stackrel{{\scriptstyle G_{\eta}}}{{\longrightarrow}}\eta_{Q}\cdot Q_{L}$
$u_{R}\stackrel{{\scriptstyle G_{\eta}}}{{\longrightarrow}}\eta_{u}\cdot u_{R},$ $d_{R}\stackrel{{\scriptstyle G_{\eta}}}{{\longrightarrow}}\eta_{d}\cdot d_{R}$
The Yukawa interactions in Eq. (32) then give the constraints:
$\eta_{Q}^{\dagger}\eta_{d}\eta_{H}=1\hskip 28.45274pt{\rm and}\hskip 28.45274pt\eta_{Q}^{\dagger}\eta_{u}\eta_{H}^{\dagger}=1.$ (34)
Now, assume ${\cal O}_{d=12}^{\text{W}}$ is allowed by this symmetry, while ${\cal O}_{d=10}$ is forbidden. This leads to the constraints:
$\eta({\cal O}_{d=12}^{\text{W}})\quad\to\quad\eta_{u}^{\dagger}(\eta_{d}^{\dagger})^{2}(\eta_{L})^{3}(\eta_{H})^{3}=1\hskip 28.45274pt{\rm and}\hskip 28.45274pt\eta({\cal O}_{d=10})\quad\to\quad(\eta_{d}^{\dagger})^{3}(\eta_{L})^{3}\eta_{H}\neq 1$ (35)
Using Eq. (34) one finds $\eta_{u}=\eta_{d}\eta_{H}^{2}$. Inserting this relation into Eq. (35) yields:
$\eta({\cal O}_{d=12}^{\text{W}})=\eta({\cal O}_{d=10}),$ (36)
which contradicts our assumption. Thus, either both operators, ${\cal O}_{d=12}^{\text{W}}$ and ${\cal O}_{d=10}$ are allowed or both are forbidden.
This symmetry argument makes it clear, why invisible neutron decay can not exist in SMEFT without nucleon decays with charged leptons in the final state, in principle. However, this kind of “proof” based on symmetries strictly speaking only guarantees that the Wilson coefficients of the operators of interest are both non-zero – it does not require them to be equal. We stress, however, even if $c_{10}$ were zero at a high scale, RGE running would generate a non-zero $c_{10}$ at the low scale – precisely because $c_{10}$ can not be protected by a symmetry.
Since the RGEs for these high-dimensional operators are not known, we can not quantify the effect. Instead, we will ask the question, how much fine-tuning one would need to accept in order to suppress decays generated from ${\cal O}_{d=10}$ to an acceptable rate. To make a quantitative statement, we therefore compare the predicted half-lives for the ${\cal O}_{d=12}^{\text{W}}$ and ${\cal O}_{d=10}$ operators with the corresponding experimental sensitivities. For ${\cal O}_{d=12}^{\text{W}}$, Eq. (2) gives our estimate. For ${\cal O}_{d=10}$, we estimate the partial width for the decay $n\to e^{-}\pi^{+}\nu\nu$ as:
$\Gamma\sim\frac{W_{0}(\pi)^{2}m_{N}^{7}\langle H^{0}\rangle^{2}}{589824\pi^{5}\Lambda^{12}}$ $\Rightarrow$ $\frac{1}{3.3\cdot 10^{23}\;{\rm yr}}\Big(\frac{13.4\;{\rm TeV}}{\Lambda}\Big)^{12}.$ (37)
Both, Eq. (2) and Eq. (37) assume the Wilson coefficient to be equal to one, and the widths scale as the Wilson coefficients squared.
[…]